hhanan:

4. Suppose the population of a town is 27,000 in 2021. The population increases at a rate of 3.7 percent every year. What will the population of the town be in 2026? Round your answer to the nearest whole number.

ZombieLover:

P= P0*(1+i/100)^t = 2700*1.04^t Where P is the population after a certain time, P0 is the initial population, t is the time since the population is P0 and i is the growing rate. To know the population after 12 years we just need to plug a 12 into the t. P=2700*1.04^12= ?

hhanan:

4322.786990 so 4323?

ZombieLover:

correct :)